OpenStudy (anonymous):

What is the equivalent of 60 miles per hour in meters per second ? ( 1 meter = 3.3 feet

OpenStudy (anonymous):

u will need this thing 1 mile = 1.6 km = 1600 mtr

OpenStudy (anonymous):

what

OpenStudy (anonymous):

u need to convert from mile/hr to meters / second

OpenStudy (anonymous):

60miles/hour =(60*1600/3600) m/sec = 160/6 m/sec = 80/3 m/sec = 26.667 m/sec

OpenStudy (anonymous):

thanks so much

OpenStudy (anonymous):

welcome :)